At work one day, Erica Franz received 45 packages. Speedy Delivery delivered three times as many as ralph's Express, while Ralph's express delivered five more than Send Quick Package Service. How many packages did each service deliver to Erica?

Found 3 solutions by rfer, stanbon, rapaljer:
Answer by rfer(16322)

s=3(x+5)=30
r=x+5=10
sq=x=5
------------
3(x+5)+(x+5)+x=45
3x+15+x+5+x=45
5x=25
x=5

Answer by stanbon(75887)

At work one day, Erica Franz received 45 packages.
s + r + q = 45
Speedy Delivery delivered three times as many as ralph's Express, while Ralph's express delivered five more than Send Quick Package Service.
s = 3r
r = s + 5
----------------------
How many packages did each service deliver to Erica?
s + r + q = 45
s - 3r + 0 = 0
s - r + 0 = -5
------------------------
I solved this system and got
unrealistic values for s,r and q.
----
Cheers,
Stan H.

Answer by rapaljer(4671)

After considerable thought, I would:

let x= number of packages from SQPS
x+5 = number of packages from RE
3(x+5) = number of packages from SD

x + x+5+3(x+5) = 45
x+x+5 + 3x + 15=45
5x + 20 = 45
5x = 25
x=5 packages from SQPS
x+5 = 10 packages from RE
3(x+5) = 30 packages from SD

If you need more help with Word Problems like these, please see my website by clicking on my tutor name "Rapaljer" anywhere in algebra.com. Look for the link on my homepage "Basic, Intermediate and College Algebra: One Step at a Time." Look for "Basic Algebra", then "Chapter 1" and select one of the topics on Word Problems. Also, see my page "MATH IN LIVING COLOR" and also my new page of videos on Factoring.

R^2

Dr. Robert J. Rapalje
Seminole Community College
Altamonte Springs, Florida